James has some $5 bills and some $20 bills. He has a total of 28 bills, and their total value is $350. Label the steps of the problem-solving process, and determine how many $5 bills James has.

Answer by addingup(3677)

Let's call the 5 x and the 20 z:
x+z= 28. If we subtract z from both sides we get:
x= 28-z Let's use this value for x in the next equation:
---------------------------------------------
5(28-z)+20z= 350
140-5z+20z= 350 Subtract 140 on both sides and add on left:
15z= 210
z= 210/15
z= 14 he has 14 $5 bills
And how many $20 bills+ 28-14
J